OpenStudy (anonymous):

show all work. Find the difference: -5(x^2 + 2x – 5) – (7x^2 – 2x – 3)

OpenStudy (anonymous):

Distribute the -5 into the first polynomial. Then distribute the -1 into the second polynomial. Then combine like terms.

OpenStudy (anonymous):

-5x^2 -10x +25 -7x^2 +4x +3

OpenStudy (anonymous):

combine like terms and that is all.

OpenStudy (anonymous):

-5x^2 - 7x^2 + 10x + 4x + 25 + 3 ?

OpenStudy (anonymous):

Yes, you put the like terms near each other. Now combine.... -5x^2 -7x^2 = -12x^2.....do that with the x-term and the constants.

OpenStudy (anonymous):

i am lost after that..

OpenStudy (anonymous):

-5x^2 - 7x^2 = -12x^2 10x+4x = 14x 25 +3 = 28 -12x^2 + 14x +28

OpenStudy (anonymous):

Sorry. You got it though.

OpenStudy (anonymous):

You wrote 28 twice...but the rest is right.

OpenStudy (anonymous):

you have 28 - 12x^2 + 14x + 28 as the final answer above both 28's supposed to be there?

OpenStudy (anonymous):

I wrote 25+3 = 28...then the next line i started the final answer. Only one 28 is needed

OpenStudy (anonymous):

oh ok i c thank you my mistake appreciate yur help

OpenStudy (anonymous):

No problem. You're welcome.